# On the Identification of Temporally Causal Representation with Instantaneous Dependence

### v3

@@ 3.3 Comparasion with Existing Methods
The application of real-world scenarios presents numerous new opportunities and challenges. Compared to existing methods of temporally causal representation learning with instantaneous effects, our approach significantly relaxes the constraints on theoretical assumptions, as shown in Table 1. TDRL (yao2022temporally) requires that there are no instantaneous effects during the transition process, which can be considered as a special case of our method. G-CaRL (morioka2023causal) and iCITRIS (lippe2023causal) require extra intervention or grouping information, which is not necessary in our method. zhang2024causal focus on non-temporal scenarios and require extra domain labels. Instead, IDOL can handle the stationary temporal case. Besides, when the instantaneous effect is dense, which is beyond the capacity of zhang2024causal, IDOL can still utilize the contextual information for better identifiability as long as the temporal transition is sparse.
@@ 4.2 Prior Estimation
To enforce the independence of noise in Equation (2), we minimize the Kullback-Leibler (KL) divergence between the posterior distribution $\prod_{i}p(z_{t,i}|\text{Pa}_{d}(z_{t,i}),\text{Pa}_{e}(z_{t,i}),{\mathbf{x}}_{t})$ and a prior $\prod_{i}p(z_{t,i}|\text{Pa}_{d}(z_{t,i}),\text{Pa}_{e}(z_{t,i}))$. This minimization implies that latent variables are mutually independent and conditioned on their historical and instantaneous parents. However, since the prior distribution may have any arbitrary density function, it is difficult to estimate such a prior. To solve this challenge, we follow chen2024caring, yao2022temporally and propose the prior networks. Specifically, we first let $r_{i}$} be an inversed function of $f$ from Equation (2) that are implemented by normalizing flow. papamakarios2021normalizing take the estimated latent variables as input to estimate the noise term $\hat{\epsilon}_{i}$, i.e. $\hat{\epsilon}_{t,i}=r_{i}(\hat{{\mathbf{z}}}_{t},\hat{{\mathbf{z}}}_{t-1})$ [^2] We use the superscript symbol to denote estimated latent variables. And each $r_{i}$ is implemented by Multi-layer Perceptron networks (MLPs). Sequentially, we devise a transformation $\kappa:=\{\hat{{\mathbf{z}}}_{t-1},\hat{{\mathbf{z}}}_{t}\}\rightarrow\{\hat{{\mathbf{z}}}_{t-1},\hat{\bm{\epsilon}}_{t}\}$, whose Jacobian can be formalized as $\mathbf{J}_{\kappa}=\begin{pmatrix}\mathbb{I}&0\\ \mathbf{J}_{d}&\mathbf{J}_{e}\end{pmatrix}$}, where $\mathbf{J}_{d}=\left(\frac{\partial r_{i}}{\partial\hat{z}_{t-1,i}}\right)$ and $\mathbf{J}_{e}=\left(\frac{\partial r_{i}}{\partial\hat{z}_{t,i}}\right)$. Hence we have Equation (9) via the change of variables formula.
@@ 5.1.1 Experimental Setup
Data Generation. We generate the simulated time series data with the fixed latent causal process as introduced in Equations (1)-(2) and Figure 1 (c). To better evaluate the proposed theoretical results, we provide six synthetic datasets from A to F with 3, 5, 8, 8, 8, 16 latent variables, respectively. Dataset D contains no instantaneous effects, which degenerate to the TDRL (yao2022temporally) setting as introduced in Figure 2 (iii). Dataset E has a dense latent causal process that violates the assumption A3(Latent Process Sparsity). All datasets except E satisfy the assumption A3. Please refer to Appendix F.1.1 for the details of data generation and evaluation metrics.
Baselines. To evaluate the effectiveness of our method, we consider the following compared methods. First, we consider the standard $\beta$-VAE (higgins2017beta) and FactorVAE (kim2018disentangling), which ignores historical information and auxiliary variables. Moreover, we consider TDRL (yao2022temporally), iVAE (khemakhem2020variational), TCL (hyvarinen2016unsupervised), PCL (hyvarinen2017nonlinear), and SlowVAE (klindt2020towards), which use temporal information but do not assume instantaneous dependency on latent processes. Finally, we consider the iCITRIS (lippe2023causal) and G-CaRL (morioka2023causal), which are devised for latent causal processes with instantaneous dependencies but require interventions or grouping of observations. As for G-CaRL, we follow the implementation description in the original paper and assign random grouping for observations. As for iCITRIS, we assign a random intervention variable for latent variables. We repeat each method over three random seeds and report the average results.
@@ 5.2.1 Experiment Setup
Datasets. To evaluate the effectiveness of our IDOL method in real-world scenarios, we conduct experiments on two human motion datasets: Human3.6M (6682899) and HumanEva-I (DBLP:journals/ijcv/SigalBB10), which record the location and orientation of local coordinate systems at each joint. We consider these datasets since the joints can be considered as latent variables and they naturally contain instantaneous dependencies. We choose several motions to conduct time series forecasting. For the Human3.6M datasets, we consider 4 motions: Gestures (Ge), Jog (J), CatchThrow (CT), and Walking (W). For HumanEva-I dataset, we consider 6 motions: Discussion (D), Greeting (Gr), Purchases (P), SittingDown (SD), Walking (W), and WalkTogether (WT). For each dataset, we select several motions and partition them into training, validation, and test sets. Please refer the Appendix F.2.1 for the dataset descriptions.
Baselines. To evaluate the effectiveness of the proposed IDOL, we consider the following state-of-the-art deep forecasting models for time series forecasting. First, we consider the conventional methods for time series forecasting including Autoformer (wu2021autoformer), TimesNet (wu2022timesnet) and MICN (wang2022micn). Moreover, we consider several latest methods for time series analysis like CARD (wang2023card), FITS (xu2024fits), and iTransformer (liu2024itransformer). Finally, we consider the TDRL (yao2022temporally). We repeat each experiment over 3 random seeds and publish the average performance. Please refer to Appendix F.2 for more experiment details.
@@ 5.2.2 Results and Discussion
Quantitative Results: Experiment results of the real-world datasets are shown in Table 3. We also conduct the Wilcoxon signed-rank test richard2021concepts on the reported accuracies, our method significantly outperforms the baselines, with a p-value threshold of 0.05. According to the experiment results, our IDOL model significantly outperforms all other baselines on most of the human motion forecasting tasks. Specifically, our method outperforms the most competitive baseline by a clear margin of 4%-34% and promotes the forecasting accuracy substantially on complex motions like the SittingDown (SD) and the Walking (W). This is because the human motion datasets contain more complex patterns described by stable causal structures, and our method can learn the latent causal process with identifiability guarantees. It is noted that our method achieves a better performance than that of TDRL, which does not consider the instantaneous dependencies among latent variables. We also find that our model achieves a comparable performance in the motion of CatchThrow (CT), this is because the size of this dataset is too small for the model to model the distribution of the observations. We further consider a more complex synthetic mixture of human motion forecasting, experiment results are shown in Appendix F.2.2. Please refer to Appendix F.2.4 for the experiment results of ablation studies. Moreover, we also investigate the proposed IDOL model on other high-dimension datasets in Appendix H.1.
@@ A.1 Identifiability of Causal Representation Learning
To achieve the causal representation (rajendran2024learning; mansouri2023object; wendong2024causal) for time series data, several researchers leverage the independent component analysis (ICA) to recover the latent variables with identification guarantees (yao2023multi; scholkopf2021toward; Liu2023CausalTriplet; gresele2020incomplete). Conventional methods assume a linear mixing function from the latent variables to the observed variables (comon1994independent; hyvarinen2013independent; lee1998independent; zhang2007kernel). To relax the linear assumption, researchers achieve the identifiability of latent variables via nonlinear ICA by using different types of assumptions like auxiliary variables or sparse generation process (zheng2022identifiability; hyvarinen1999nonlinear; hyvarinen2023identifiability; khemakhem2020ice; li2023identifying). As for the methods that employ the auxiliary variables, khemakhem2020variational first achieve the identifiability by assuming the latent sources with exponential family and introducing auxiliary variables e.g., domain indexes, time indexes, and class labels (khemakhem2020variational; hyvarinen2016unsupervised; hyvarinen2017nonlinear; hyvarinen2019nonlinear). To further relax the exponential family assumption, kong2022partial; xie2022multi; kong2023identification; yan2023counterfactual; xie2022multi achieve the component-wise identification results for nonlinear ICA with a $2n+1$ number of auxiliary variables for $n$ latent variables. Recently, li2024subspace further relax to $n+1$ number of auxiliary variables and achieve subspace identifiability. To seek identifiability in an unsupervised manner, researchers employ the assumption of structural sparsity to achieve identifiability (ng2024identifiability; lachapelle2022disentanglement; zheng2022identifiability; xu2024sparsity). Specifically, lachapelle2023synergies; lachapelle2022partial proposed mechanism sparsity regularization as an inductive bias to identify the causal latent factors and achieve the identifiability on several scenarios like multi-task learning. They further show the identifiability results up to the consistency relationship (lachapelle2024nonparametric), which allows the partial disentanglement of latent variables. Recently, (zhang2024causal) use the sparse structures of latent variables to achieve identifiability under distribution shift. Different from these methods that assume sparsity exists in the generation process between the latent source and the observation, we assume the sparse latent process, where sparsity exists in the transition of latent variables and is common in real-world time series data.
@@ A.2 Nonlinear ICA for Time Series Data
Our work is also related to the temporally causal representation. Existing methods for temporally causal representation (yan2023counterfactual; huang2023latent; halva2020hidden; lippe2022citris) usually rely on the conditional independence assumption, where the latent variables are mutually independent conditional on their time-delayed parents. Specifically, hyvarinen2016unsupervised leverage the independent sources principle and the variability of data segments to achieve identifiability on nonstationary time series data. They further use permutation-based contrastive (hyvarinen2017nonlinear) to achieve identifiability on stationary time series data. The recent advancements in temporally causal representation include LEAP (yao2021learning), TDRL (yao2022temporally), which leverage the properties of independent noises and variability historical information. However, this assumption is too strong to meet in real-world scenarios since instantaneous dependencies are common in real-world scenarios like human motion. To solve this problem, researchers introduce further assumptions. For example, lippe2023causal propose the i-Citris, which demonstrates identifiability by assuming interventions on latent variables. In addition, morioka2023causal propose G-CaRL, which yields identifiability with the assumption of the grouping of observational variables. However, these assumptions are still hard to meet in practice. Hence we propose a more relaxed assumption, transition sparsity, to achieve identifiability under time series data with instantaneous
@@ A.3 Causal Discovery with Latent Variables
Several studies are proposed to discover causally related latent variables (chen2024identification; lachapelle2024additive; monti2020causal; zhang2012identifiability; tashiro2012estimation). Specifically, (huang2022latent; kong2024identification) leverage the vanishing Tetrad conditions or rank constraints to identify latent variables in linear-Gaussian models. xie2022identification; cai2019triad further draw upon non-Gaussianity in their analysis for linear, non-Gaussian scenarios. Other methods aim to reconstruct the hierarchical structures of the latent variables, like xie2022identification; huang2022latent. However, these methods usually use the linear assumption and can hardly handle the real-world time series data with complex nonlinear relationships.
@@ A.4 Instantaneous Dependencey of Time Series Data
When the discrete time series data is sampled in a low-frequency or low-resolution manner, instantaneous dependence (adak1996time; koutlis2019identification; gersch1985modeling; jamaludeen2022discovering; wu2023hierarchical) occurs where variables at each time step are not independent given their historical observations. swanson1997impulse first discuss the instantaneous dependence from the perspective of noise. Lots of works investigate instantaneous dependency from the causal view. Specifically, hyvarinen2010estimation estimate both instantaneous and lagged effects via a non-Gaussian model. gong2015discovering; gong2017causal discover causal structure with instantaneous dependency from subsampled and aggregated data. Recently, zhu2023beware consider instantaneous dependency in reinforcement learning.
@@ A.5 Time Series Forecasting
Time series forecasting (hyndman2018forecasting; box1970distribution) is one of the most popular research problems in recent years. Recently, the deep-learning-based methods have made great progress, which can be categorized according to different types of neural architectures like the RNN-based methods (hochreiter1997long; lai2018modeling; salinas2020deepar), CNN-based models (bai2018empirical; wang2022micn; wu2022timesnet), and the methods based on state-space model (gu2022parameterization; gu2021combining; gu2021efficiently). Recently, Transformer-based methods (zhou2021informer; wu2021autoformer; nie2022time) further push the development of time series forecasting. However, these methods seldom consider the instantaneous dependencies of time series data.

### v4

@@ 3.3 Comparasion with Existing Methods
The application of real-world scenarios presents numerous new opportunities and challenges. Compared to existing methods of temporally causal representation learning with instantaneous effects, our approach significantly relaxes the constraints on theoretical assumptions, as shown in Table 1. TDRL (Yao et al., 2022) requires that there are no instantaneous effects during the transition process, which can be considered as a special case of our method. G-CaRL (Morioka and Hyvärinen, 2023) and iCITRIS (Lippe et al., 2023) require extra intervention or grouping information, which is not necessary in our method. Zhang et al. (2024) focus on non-temporal scenarios and require extra domain labels. Instead, IDOL can handle the stationary temporal case. Besides, when the instantaneous effect is dense, which is beyond the capacity of Zhang et al. (2024), IDOL can still utilize the contextual information for better identifiability as long as the temporal transition is sparse.
@@ 4.2 Prior Estimation
To enforce the independence of noise in Equation (2), we minimize the Kullback-Leibler (KL) divergence between the posterior distribution $\prod_{i}p(z_{t,i}|\text{Pa}_{d}(z_{t,i}),\text{Pa}_{e}(z_{t,i}),{\mathbf{x}}_{t})$ and a prior $\prod_{i}p(z_{t,i}|\text{Pa}_{d}(z_{t,i}),\text{Pa}_{e}(z_{t,i}))$. This minimization implies that latent variables are mutually independent and conditioned on their historical and instantaneous parents. However, since the prior distribution may have any arbitrary density function, it is difficult to estimate such a prior. To solve this challenge, we follow Chen et al. (2024a), Yao et al. (2022) and propose the prior networks. Specifically, we first let $r_{i}$} be an inversed function of $f$ from Equation (2) that are implemented by normalizing flow. Papamakarios et al. (2021) take the estimated latent variables as input to estimate the noise term $\hat{\epsilon}_{i}$, i.e. $\hat{\epsilon}_{t,i}=r_{i}(\hat{{\mathbf{z}}}_{t},\hat{{\mathbf{z}}}_{t-1})$ [^2] We use the superscript symbol to denote estimated latent variables. And each $r_{i}$ is implemented by Multi-layer Perceptron networks (MLPs). Sequentially, we devise a transformation $\kappa:=\{\hat{{\mathbf{z}}}_{t-1},\hat{{\mathbf{z}}}_{t}\}\rightarrow\{\hat{{\mathbf{z}}}_{t-1},\hat{\bm{\epsilon}}_{t}\}$, whose Jacobian can be formalized as $\mathbf{J}_{\kappa}=\begin{pmatrix}\mathbb{I}&0\\ \mathbf{J}_{d}&\mathbf{J}_{e}\end{pmatrix}$}, where $\mathbf{J}_{d}=\left(\frac{\partial r_{i}}{\partial\hat{z}_{t-1,i}}\right)$ and $\mathbf{J}_{e}=\left(\frac{\partial r_{i}}{\partial\hat{z}_{t,i}}\right)$. Hence we have Equation (9) via the change of variables formula.
@@ 5.1.1 Experimental Setup
Data Generation. We generate the simulated time series data with the fixed latent causal process as introduced in Equations (1)-(2) and Figure 1 (c). To better evaluate the proposed theoretical results, we provide six synthetic datasets from A to F with 3, 5, 8, 8, 8, 16 latent variables, respectively. Dataset D contains no instantaneous effects, which degenerate to the TDRL (Yao et al., 2022) setting as introduced in Figure 2 (iii). Dataset E has a dense latent causal process that violates the assumption A3(Latent Process Sparsity). All datasets except E satisfy the assumption A3. Please refer to Appendix F.1.1 for the details of data generation and evaluation metrics.
Baselines. To evaluate the effectiveness of our method, we consider the following compared methods. First, we consider the standard $\beta$-VAE (Higgins et al., 2017) and FactorVAE (Kim and Mnih, 2018), which ignores historical information and auxiliary variables. Moreover, we consider TDRL (Yao et al., 2022), iVAE (Khemakhem et al., 2020a), TCL (Hyvarinen and Morioka, 2016), PCL (Hyvarinen and Morioka, 2017), and SlowVAE (Klindt et al., 2020), which use temporal information but do not assume instantaneous dependency on latent processes. Finally, we consider the iCITRIS (Lippe et al., 2023) and G-CaRL (Morioka and Hyvärinen, 2023), which are devised for latent causal processes with instantaneous dependencies but require interventions or grouping of observations. As for G-CaRL, we follow the implementation description in the original paper and assign random grouping for observations. As for iCITRIS, we assign a random intervention variable for latent variables. We repeat each method over three random seeds and report the average results.
@@ 5.2.1 Experiment Setup
Datasets. To evaluate the effectiveness of our IDOL method in real-world scenarios, we conduct experiments on two human motion datasets: Human3.6M (Ionescu et al., 2014) and HumanEva-I (Sigal et al., 2010), which record the location and orientation of local coordinate systems at each joint. We consider these datasets since the joints can be considered as latent variables and they naturally contain instantaneous dependencies. We choose several motions to conduct time series forecasting. For the Human3.6M datasets, we consider 4 motions: Gestures (Ge), Jog (J), CatchThrow (CT), and Walking (W). For HumanEva-I dataset, we consider 6 motions: Discussion (D), Greeting (Gr), Purchases (P), SittingDown (SD), Walking (W), and WalkTogether (WT). For each dataset, we select several motions and partition them into training, validation, and test sets. Please refer the Appendix F.2.1 for the dataset descriptions.
Baselines. To evaluate the effectiveness of the proposed IDOL, we consider the following state-of-the-art deep forecasting models for time series forecasting. First, we consider the conventional methods for time series forecasting including Autoformer (Wu et al., 2021), TimesNet (Wu et al., 2022) and MICN (Wang et al., 2022). Moreover, we consider several latest methods for time series analysis like CARD (Wang et al., 2023), FITS (Xu et al., 2024b), and iTransformer (Liu et al., 2024). Finally, we consider the TDRL (Yao et al., 2022). We repeat each experiment over 3 random seeds and publish the average performance. Please refer to Appendix F.2 for more experiment details.
@@ 5.2.2 Results and Discussion
Quantitative Results: Experiment results of the real-world datasets are shown in Table 3. We also conduct the Wilcoxon signed-rank test Richard (2021) on the reported accuracies, our method significantly outperforms the baselines, with a p-value threshold of 0.05. According to the experiment results, our IDOL model significantly outperforms all other baselines on most of the human motion forecasting tasks. Specifically, our method outperforms the most competitive baseline by a clear margin of 4%-34% and promotes the forecasting accuracy substantially on complex motions like the SittingDown (SD) and the Walking (W). This is because the human motion datasets contain more complex patterns described by stable causal structures, and our method can learn the latent causal process with identifiability guarantees. It is noted that our method achieves a better performance than that of TDRL, which does not consider the instantaneous dependencies among latent variables. We also find that our model achieves a comparable performance in the motion of CatchThrow (CT), this is because the size of this dataset is too small for the model to model the distribution of the observations. We further consider a more complex synthetic mixture of human motion forecasting, experiment results are shown in Appendix F.2.2. Please refer to Appendix F.2.4 for the experiment results of ablation studies. Moreover, we also investigate the proposed IDOL model on other high-dimension datasets in Appendix H.1.
@@ A.1 Identifiability of Causal Representation Learning
To achieve the causal representation (Rajendran et al., 2024; Mansouri et al., 2023; Wendong et al., 2024) for time series data, several researchers leverage the independent component analysis (ICA) to recover the latent variables with identification guarantees (Yao et al., 2023; Schölkopf et al., 2021; Liu et al., 2023; Gresele et al., 2020). Conventional methods assume a linear mixing function from the latent variables to the observed variables (Comon, 1994; Hyvärinen, 2013; Lee and Lee, 1998; Zhang and Chan, 2007). To relax the linear assumption, researchers achieve the identifiability of latent variables via nonlinear ICA by using different types of assumptions like auxiliary variables or sparse generation process (Zheng et al., 2022; Hyvärinen and Pajunen, 1999; Hyvärinen et al., 2023; Khemakhem et al., 2020b; Li et al., 2023c). As for the methods that employ the auxiliary variables, Khemakhem et al. (2020a) first achieve the identifiability by assuming the latent sources with exponential family and introducing auxiliary variables e.g., domain indexes, time indexes, and class labels (Khemakhem et al., 2020a; Hyvarinen and Morioka, 2016; 2017; Hyvarinen et al., 2019). To further relax the exponential family assumption, Kong et al. (2022); Xie et al. (2022b); Kong et al. (2023); Yan et al. (2023); Xie et al. (2022b) achieve the component-wise identification results for nonlinear ICA with a $2n+1$ number of auxiliary variables for $n$ latent variables. Recently, Li et al. (2024) further relax to $n+1$ number of auxiliary variables and achieve subspace identifiability. To seek identifiability in an unsupervised manner, researchers employ the assumption of structural sparsity to achieve identifiability (Ng et al., 2024; Lachapelle et al., 2022; Zheng et al., 2022; Xu et al., 2024a). Specifically, Lachapelle et al. (2023); Lachapelle and Lacoste-Julien (2022) proposed mechanism sparsity regularization as an inductive bias to identify the causal latent factors and achieve the identifiability on several scenarios like multi-task learning. They further show the identifiability results up to the consistency relationship (Lachapelle et al., 2024a), which allows the partial disentanglement of latent variables. Recently, (Zhang et al., 2024) use the sparse structures of latent variables to achieve identifiability under distribution shift. Different from these methods that assume sparsity exists in the generation process between the latent source and the observation, we assume the sparse latent process, where sparsity exists in the transition of latent variables and is common in real-world time series data.
@@ A.2 Nonlinear ICA for Time Series Data
Our work is also related to the temporally causal representation. Existing methods for temporally causal representation (Yan et al., 2023; Huang et al., 2023; H"alv"a and Hyvarinen, 2020; Lippe et al., 2022) usually rely on the conditional independence assumption, where the latent variables are mutually independent conditional on their time-delayed parents. Specifically, Hyvarinen and Morioka (2016) leverage the independent sources principle and the variability of data segments to achieve identifiability on nonstationary time series data. They further use permutation-based contrastive (Hyvarinen and Morioka, 2017) to achieve identifiability on stationary time series data. The recent advancements in temporally causal representation include LEAP (Yao et al., 2021), TDRL (Yao et al., 2022), which leverage the properties of independent noises and variability historical information. However, this assumption is too strong to meet in real-world scenarios since instantaneous dependencies are common in real-world scenarios like human motion. To solve this problem, researchers introduce further assumptions. For example, Lippe et al. (2023) propose the i-Citris, which demonstrates identifiability by assuming interventions on latent variables. In addition, Morioka and Hyvärinen (2023) propose G-CaRL, which yields identifiability with the assumption of the grouping of observational variables. However, these assumptions are still hard to meet in practice. Hence we propose a more relaxed assumption, transition sparsity, to achieve identifiability under time series data with instantaneous
@@ A.3 Causal Discovery with Latent Variables
Several studies are proposed to discover causally related latent variables (Chen et al., 2024b; Lachapelle et al., 2024b; Monti et al., 2020; Zhang and Hyvarinen, 2012; Tashiro et al., 2012). Specifically, (Huang et al., 2022; Kong et al., 2024) leverage the vanishing Tetrad conditions or rank constraints to identify latent variables in linear-Gaussian models. Xie et al. (2022a); Cai et al. (2019) further draw upon non-Gaussianity in their analysis for linear, non-Gaussian scenarios. Other methods aim to reconstruct the hierarchical structures of the latent variables, like Xie et al. (2022a); Huang et al. (2022). However, these methods usually use the linear assumption and can hardly handle the real-world time series data with complex nonlinear relationships.
@@ A.4 Instantaneous Dependencey of Time Series Data
When the discrete time series data is sampled in a low-frequency or low-resolution manner, instantaneous dependence (Adak, 1996; Koutlis et al., 2019; Gersch, 1985; Jamaludeen et al., 2022; Wu et al., 2023a) occurs where variables at each time step are not independent given their historical observations. Swanson and Granger (1997) first discuss the instantaneous dependence from the perspective of noise. Lots of works investigate instantaneous dependency from the causal view. Specifically, Hyvärinen et al. (2010) estimate both instantaneous and lagged effects via a non-Gaussian model. Gong et al. (2015; 2017) discover causal structure with instantaneous dependency from subsampled and aggregated data. Recently, Zhu et al. (2023) consider instantaneous dependency in reinforcement learning.
@@ A.5 Time Series Forecasting
Time series forecasting (Hyndman and Athanasopoulos, 2018; Box and Pierce, 1970) is one of the most popular research problems in recent years. Recently, the deep-learning-based methods have made great progress, which can be categorized according to different types of neural architectures like the RNN-based methods (Hochreiter and Schmidhuber, 1997; Lai et al., 2018; Salinas et al., 2020), CNN-based models (Bai et al., 2018; Wang et al., 2022; Wu et al., 2022), and the methods based on state-space model (Gu et al., 2022; 2021b; 2021a). Recently, Transformer-based methods (Zhou et al., 2021; Wu et al., 2021; Nie et al., 2022) further push the development of time series forecasting. However, these methods seldom consider the instantaneous dependencies of time series data.